OpenStudy (anonymous):

12.25 = 0.5q + 3.75 What is the answer

OpenStudy (lucaz):

subtract 3.75 from both sides and divide both sides by 0.5 8.5=0.5q q=8.5/0.5=17

OpenStudy (anonymous):

what do you mean

OpenStudy (unklerhaukus):

12.25 = 0.5q + 3.75 12.25 -3.75 = 0.5q + 3.75 -3.75 12.25 -3.75 = 0.5q (12.25 -3.75)/0.5 = 0.5q/0.5 (12.25 -3.75)/0.5 = q

OpenStudy (anonymous):

Is the answer 4.75

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

14a + 10a = 72 what is the answer

OpenStudy (unklerhaukus):

14a + 10a = 72 (14 + 10)a = 72 a = 72/(14+10)